OpenStudy (anonymous):

would some body help with 2 math probs #1 estimate the diameter of the circle with the circumference of 80 feet #2 Suzanne wants to put a fence around her square garden if the garden covers an area of 100 ft how many feet of fencing dose she need

OpenStudy (anonymous):

#1 Using the circumference of the circle formula.. C = πd Given that C = 80.. 80 = πd d = 80/π ft ≈ 25.4 ft

OpenStudy (anonymous):

thanks what about #2

OpenStudy (anonymous):

#2 It's 40 the explanation is long but I can kind of describe it if you have 100 square feet the area of a square is equal to x^2 = area, so x^2 = 100 you must then take the square root of each side to get rid of the square and the square root of 100 is 10 or -10 but you can't have -10 because it's distance in fence so that can't be an answer so you are left with 10. The perimeter of a fence is equal to 4x so 4(10) = 40 then put your unit back in and you have 40 feet of fencing.

OpenStudy (anonymous):

would you help me with 2 more questons Plz

OpenStudy (anonymous):

okie i will try

OpenStudy (anonymous):

the diameter of a tire is 2.5 ft use this mesuerment to answer a and b a. find the circunfrince of the tire b. about how many times will the tier have to rotate to travel 1mile = 5280 ft